Summary of the GMVA Session, Sep. 30 - Oct 4, 2021 Ver.3, TPK, Oct. 19, 2021 ************************************************** Stations participating: Europe: Ef, On, Mh, Ys, Pv, NOEMA(phased) US: 8xVLBA at 3mm, 10xVLBA at 7mm GBT at 3mm KVN: all =============================================================================================== European Fringetest: f212a - good fringes found for Ef, On, Mh, Pv, Noema cross hand SNR is a bit high for Mh fringes to NOEMA are particularly strong in the beginning NOEMA showed fringes only in one polarisation (RCP) due to time sync problem in PolyFiX. Details here: https://deki.mpifr-bonn.mpg.de/GMVA/GMVA_correlations/2021_Session_II_(c212)/F212A =============================================================================================== Overall Summary: in Effelsberg and northern Europe the weather was rather poor with rain and wind, affecting Ef+On+Mh. PV had better weather. Noema was good in the first half of the session, but degrading towards the end. For a certain fraction of the session less than 10 antennas were phased. A factor of two reduction in the slew rates for GBT was anounced late and may have affect source switching. The VLBA is affected at some stations by a missing 2nd polarisation (PT, SC, HN). =============================================================================================== Detailed Notes from Stations: Effelsberg: ------------ Observations went without know technical problems. The weather conditions were moderate at the 1st half of of the GMVA session (cloudy) and substially degrades in the 2nd half (rain, high wind). Pointing was often done only with the P10mm receiver; turning the receiver from 3mm to 10mm takes several minutes, so pointing cadence was lower than usual. The alignement between the 3mm and 10mm receiver is thought to be within a few arcseconds (beam@3mm: 10.5"). Below we list downtimes, longer than a single scan: f212a - all scans observed c212a - the following scans not observed 273-23h00 -> 23h30 274-13h00 -> 13h52 274-20h40 -> 20h51 c212b 275-19h30 -> 276-03h37 stop due to high wind 276-09h15 -> until end of c212b (276-13h37) c212c 276-14h00 -> 276-14h52 stop to high wind 276-15h30 -> 275-15h52 276-17h20 -> 276-17h31 276-21h40 -> 276-21h52 277-00h00 -> end of c212c was observed, but the weather was extremly bad (wind, rain, thick clouds) Onsala: -------- The Onsala 20m telescope successfully participated in the GMVA session 2021/2 with the 3mm receiver. The fringes were found during the FTP-Fringe tests of F212A. Before and after the observations, we checked the test signal to verify the LO correctness. The 3-mm pointing model has a reasonable accuracy (Delta AZ/EL < 5 arcsec). In the log file, I did not see any report of the data packet loss. While, the local weather was poor during the entire session. It was cloudy and raining almost all the time. The poor weather caused a humidity of ~100%. The opacity corrected Tsys* reached ~20000 K for a certain scans. Metsahovi: ----------- We did not detect any technical issues during the session and we participated to all scheduled observations (no missed scans etc.) The weather conditions were not favorable either this time. f212a: cloudy c212a: cloudy and rain c212b: cloudy c212c: cloudy, and last 10 hours rather nice weather conditions Yebes: ------- no info yet Pico Veleta: ------------ The weather was good in general. We had strong anomalous refraction in the afternoons (~12:00-18:00 UT), but very stable atmosphere outside that time range. Opacity varied along the week (0.1< tau225 < 0.6), but for 3mm it did not affect significantly the Tsys*, except for some short periods of bad weather (fog and snow during c212c). For the individual schedules: f212a - Done 100% under good weather c212a - Done 96% under good weather c212b - Done 95% under good weather. Note that between 03:00 - 09:00 UTC we received scans in the .vex but were not scheduled for VLBI observations. c212c - Done 69%, under a mix of poor/bad weather (first part) and good weather (second part) NOEMA: ------- 1) We had one technical issue, that the time server had disappeared for one of the Polyfix unit we were using for this session (unit 6), explaining the first lack of correlation in LCP. This affects all scans of f212a and all scans of c212a until the first good scan at 09h00 UT (on 1156+295). Before our gap in the night from Thursday to Saturday, files had the same size for both polar, but not on Friday from 06h20 on until 09h00. Dominique thinks that this only a integer offset in the second stored in the header (of order 10s) so that fringes could eventually be found. 2) We had a bug resulting in using only 3 antennas in the adding mask on Friday from 16h00 to 20h20 UT (c212a). Rest of the time, we had between 6 and 8 antennas in the mask (+1 comparison antenna). 3) We could not phase on 1915+105 (c212a) which was too weak (it has apparently 3 mJy flux density at 3mm). So: - first scan (17h25-17h31), we had active phasing running, giving rubbish phasing. - for all subsequent scans on 1915+105, we manually deactivated active phasing (and reactivated it on the quasars), EXCEPT for scan 21h36-21h51 where the deactivation was missed. Not sure however if and how the passive phasing holded for 6 minutes (and we have no measure of phasing efficiency when off). (but could phase correctly on SS433 with 5s integration for phasing). 4) We had to stop due to first wind, then storm and rain from Sunday 03h30 UT on until the end of the session (from c212b scan 03h30 on OJ287 to end of c212b and full c212c). We kept recording and loaded c212c schedule nonetheless. 5) Below is a list for all scans that had problems (excluding problems mentioned above): - f212a, scan 12h00-12h05 (OJ287): no 2-bit calib before start. - f212a, scan 12h15-12h20 (OJ287): data acquisition has been killed by mistake. - f212a, scan 13h45-13h50 (3C345): pb acquisitions. - c212a, scans 20h20-20h23, 20h25-20h31, 20h40-20h43 (1749+096, 1915+105, 1749+096): pb acquisition unit 6. - c212a, scans 08h30-08h37, 08h45-08h52, 09h00-09h7 (1156+295): pb acquisition unit 6 (pb of time server was being patched). - c212a, scan 11h15-11h22: hot load in front of the receiver (at least part of the scan). - c212a, scan 12h50-12h55: missed scan (taking an antenna out). - c212b, scans 10h30-10h37, 10h56-10h52, 11h00-11h07: missed scans (alarm not reporting). - c212b, scan 14h45-14h52: missed scan (online computer problem). KVN: ----- no info yet =============================================================================================== VLBA: ------ Inspection of Tsys values from VLBA (used as proxy of performance): For all parts of the session: Tsys is close to 1000K for several stations for some 50% of the data, more often in LCP than in RCP. It is unclear if this problem comes from a ploting script or is physically real. c212a: BR 3mm ok, 7mm ok FD 3mm LCP 2x higher than RCP ; 7mm ok HN no 3mm receiver, 7mm LCP missing KP 3mm ok, 7mm ok LA 3mm ok, 7mm ok MK 3mm ok, 7mm ok NL 3mm ok, 7mm ok OV 3mm ok, 7mm ok PT 3mm RCP ok, LCP missing, 7mm ok SC no 3mm receiver, 7mm LCP missing GBT no Tsys data c212b: BR 3mm ok 7mm ok FD 3mm LCP 1.3x higher than RCP ; 7mm ok HN no 3mm receiver, 7mm LCP missing KP 3mm ok, 7mm ok LA 3mm ok, 7mm ok MK 3mm ok, 7mm ok NL 3mm ok, 7mm ok NL 3mm ok, 7mm ok PT 3mm RCP ok, LCP missing, 7mm ok SC no 3mm receiver, 7mm LCP missing GBT should have observed, but no files on /astronomy/oct21/c212b/ c212c: BR 3mm ok, 7mm ok FD 3mm LCP 1.5x higher than RCP ; 7mm ok KP 3mm ok, 7mm ok LA 3mm ok, 7mm ok MK 3mm ok, 7mm ok NL 3mm ok, 7mm ok PT 3mm RCP ok, LCP 3xhigher Tsys, 7mm ok SC no 3mm receiver, 7mm LCP has only few entries (~50% missing) =============================================================================================== VLBA Observers Log VERY LONG BASELINE ARRAY OBSERVING LOG -------------------------------------- Project: C212A Observer: Krichbaum, T. Project type: VLBA Obs filename: c212a.vex Date/Day: 2021SEP30/273 Ants Scheduled: SC HN NL FD LA PT KP OV BR MK GB =UT-Time===Comment===============================================MF#===%AD==AMD= Operator is Kenneth Gibson 2052 Begin at NL FD LA PT KP OV BR 2110 Begin at SC HN 2147 New operator is Carolyn Morris 2248 Begin at MK 0000 Date is 2021OCT01/274. 0026 %MK raining. 0410 Begin at GB. 0410-0707 *GB not observing, telescope related software SW 100 177 problem on GBT side. 0510-0455 %SC 7mm cryo is warm, but cooling CRYO 0538 New Operator: Paul Padilla 0548-0550 *KP autostowed, power glitch COM 100 2 0709-0712 *KP autostowed, power glitch COM 100 3 1039 %PT raining. 1351 New operator is Betty Ragan 1408 %BR heavy dew on dish. 1806-1953 *HN, MK removed for regular USNO observation. SCH 200 214 1933 %SC raining. 2148 New operator is Carolyn Morris 2330-0455 *SC 7mm warm again, very slowly cooling but tsys bad CRYO 100 325 0450 End at GB. 0455 End. Downtime Summary: Total downtime : 721 min Percentage downtime of observing: 14.9% Average downtime per hour : 9.0 min Total scheduled observing time (# Antennas): 4830 min (10) VERY LONG BASELINE ARRAY OBSERVING LOG -------------------------------------- Project: C212B Observer: Krichbaum, T. Project type: VLBI Obs filename: c212b.vex Date/Day: 2021OCT02/275 Ants Scheduled: SC HN NL FD LA PT KP OV BR MK GB =UT-Time===Comment===============================================MF#===%AD==AMD= Operator: Carolyn Morris 0459 Begin at NL,FD,LA,PT,KP,OV,BR 0545 New Operator: Paul Padilla 0638 MK joins 0721-0833 *MK out for scheduled Q series USNO SCH 100 72 observation 1307-1312 *MK removed temporarily. I had noticed that HUERR 100 5 there is a gap in MK's schedule. It seems a previously queued project that interrupted this one was removed, but this project was not correctly re-queued afterwards. I am doing that now. 1400 New operator is Alan Kerr 1912 Begin at GB 1912-2023 *GB timing problems. was not caught earlier TIME 100 71 because the alerts would clear at the end of every scan. 2036 SC,HN begin. 0000 Date is 2021OCT03/276. 0023-0027 *HN autostow. power glitch? repointed. COM 100 4 Operator: Jessica King 0207-0217 *HN source was directly at zenith. Antenna was OTHER 100 10 unable to track and autostowed itself. 0705-0706 *HN autostowed again, not sure why this time. COM 100 1 Possibly a power glitch. 0709 SC end. 0735 GB end. 0809 HN end. 1345 New operator is Alan Kerr 1408 End. Downtime Summary: Total downtime : 163 min Percentage downtime of observing: 0.8% Average downtime per hour : 0.5 min Total scheduled observing time (# Antennas): 19890 min (10) VERY LONG BASELINE ARRAY OBSERVING LOG -------------------------------------- Project: C212C Observer: Krichbaum, T. Project type: VLBI Obs filename: c212c.vex Date/Day: 2021OCT03/276 Ants Scheduled: SC HN NL FD LA PT KP OV BR MK =UT-Time===Comment===============================================MF#===%AD==AMD= Operator is Alan Kerr 1400 Begin at non-VLBA sites. 2353 Begin at NL,FD,LA,PT,KP,OV,BR 2359 Begin at SC,HN 0000 Date is 2021OCT04/277. 0107 %HN rain 0150 New operator: Jessica King 0450 MK begin. 1345 New operator is Alan Kerr 1359 End. Downtime Summary: Total downtime : 0 min Percentage downtime of observing: 0.0% Average downtime per hour : 0.0 min Total scheduled observing time (# Antennas): 14390 min (10) Notes: * = Entries where data was affected. % = Entries where data may have been affected. & = Entries where the site tech was called out. WEA = Weather entries. MF# = Maintenance form or major downtime category associated with a problem. %AD = The percentage of an antenna affected by a problem. AMD = Total antenna-minutes downtime for a problem. Tsys = System Temperature (TP/SP x Tcal/2) ACU = Antenna Control Unit FRM = Focus/Rotation Mount RFI = Radio Frequency Interference VME = Site control computer CB = Circuit Breaker vclock = Program that compares site clock time to a standard.